# Tackling Cyberattacks through AI-based Reactive Systems: A Holistic Review and Future Vision

### v1

@@ Abstract
There is no denying that the use of Information Technology (IT) is undergoing exponential growth in today’s world. This digital transformation has also given rise to a multitude of security challenges, notably in the realm of cybercrime. In response to these growing threats, public and private sectors have prioritized the strengthening of IT security measures. In light of the growing security concern, Artificial Intelligence (AI) has gained prominence within the cybersecurity landscape. This paper presents a comprehensive survey of recent advancements in AI-driven threat response systems. To the best of our knowledge, the most recent survey covering the AI reaction domain was conducted in 2017. Since then, considerable literature has been published and therefore it is worth reviewing it. By means of several shared features, each of the studies is compared on a common ground. Through an analysis of the research papers conducted on a standardized basis, this survey aims to unravel the complexities and opportunities of integrating AI into cyber defense. The conclusions drawn from this collective analysis provide a comprehensive snapshot of the evolving landscape at the intersection of AI and cybersecurity. This landscape underscores the growing significance of not only anticipating and detecting threats but also responding to them effectively. Additionally, from these reviews, various research challenges for the future are presented. These challenges serve as a roadmap for researchers and practitioners in the field of AI-integrated reactive strategies.
@@
Artificial Intelligence, Reaction, Countermeasure, Survey
@@ I Introduction
The conjunction between AI and cybersecurity reaction systems can bring a powerful combination that amplifies the effectiveness of cyber defense mechanisms. For example, where IRS focuses on real-time threat monitoring, analysis, and incident response, AI increases these capabilities by taking advantage of advanced analytics, ML, and predictive models. In this sense, the collaboration enables the identification of sophisticated attack patterns, the prediction of unknown future threats, as well as the generation of actionable perspectives to inform decision-making processes [33].
Some researchers have recognized the significance of incorporating AI techniques into cyber defense mechanisms [34]. Furthermore, numerous papers have proposed a variety of approaches to develop more effective countermeasure selection mechanisms that use AI. Therefore, it is crucial to review and extend the findings obtained from the studies by integrating the latest advances in AI and capturing the knowledge gained from subsequent research efforts. This is why it is important to review the state of the art of threat response systems using AI so that all developments or approaches in the field can be brought into a unified vision.
@@
There have been previous studies [37, 38, 39, 40] that unify previously conducted work investigating the automatability of countermeasure selection and reaction strategies. However, due to the rapid pace of advancements in the field of reaction mechanisms using AI, there is a need for an up-to-date study to consolidate existing knowledge, establish benchmarks, and compare different modern approaches. As far as we are aware, the latest survey addressing the field of AI reaction was carried out in 2017. Therefore, by conducting a comprehensive study, a holistic understanding of the current state of the art can be gained, research gaps can be identified, and valuable information can be provided for researchers, practitioners, and organizations working to improve cybersecurity defenses through AI-based reaction systems. For this reason, the objective of this paper is to conduct a comprehensive survey and analysis of AI-based reaction systems. To do so, the most remarkable 22 papers (to the best of our knowledge) from the last 5 years have been analysed, comparing them in an equivalent basis across 5 selected features that have been considered relevant to the world of reaction. Moreover, the paper aims to individually examine and impartially compare these works. In doing so, it is intended to highlight emerging challenges and outline possible research directions for future research in the field of AI-integrated reactive strategies. Furthermore, the paper aims to locate relevant related studies that have the potential to improve the effectiveness of such systems in future efforts.
@@ II Background
For the sake of the reader, the main objective of this section is to cover several key concepts that form the basis of AI-powered threat response systems. Particularly, understanding these concepts is essential to understanding the context and importance of this research ecosystem. As discussed previously, the increase in the number of cyberattacks is tangible. This is why systems such as IDS, which are dedicated to detecting threats, are used. And, it is because of the inability of IDS to react that they are combined with response systems such as IPS or IRS.
Previously, cyberattack detection and prevention systems heavily relied on static techniques such as rule-based analysis [41, 42] or approaches built upon hash-based analysis [43] or the normal behavior of applications [44]. However, with the advancement of technology and the increasing sophistication of cyber threats, static detection has become inefficient [45]. Therefore, there is a growing need for more advanced and dynamic approaches to effectively safeguard against evolving attack vectors, that can cover modern attacks.
@@
For a better comprehension understanding of each of these systems (IDS, IPS and IRS), Figure 1 has been designed. In this figure, the different components are graphically displayed. The scene is represented in some steps, that correspond to the numbers associated in the figure: 1) The scenario describes an attack from a malicious actor where a device is infected. 2) After infection, the IDS would be in charge of detecting the malware on the infected device. 3) Once identified the malware, it would send an alarm to the IRS. In this case, the IRS would be in charge of the incident response. 4) After analyzing the alert it receives from the IPS, the IRS would then select the most effective countermeasure and implement it. 5) On the other hand, an IPS is constantly analyzing the network and proactively generating security measures.
Figure 1: Prevention, response and detection AI-based systems scenario
@@
Technologies such as AI can offer a new approach that leverages the latest advancements to create highly efficient systems[49, 50]. These new developments have the potential to enhance on many levels, for instance in the fields of medicine [51], robotics [52], finance [53], and transport [54], among others. When it comes to cyber defense, the deployment of these advances can help for example in improving the detection, prevention, and response systems discussed above. For instance, one main advantage of AI in cybersecurity is its ability to analyze large volumes of data and identify patterns and anomalies that may be indicative of a cyberattack [55]. For example, algorithms using ML can be trained on large amounts of datasets to recognize different patterns of known attacks and detect previously unseen threats, such as “zero-days” vulnerabilities [56]. As already said, traditional approaches to security rely on static, rule-based systems that are unable to handle the increasing complexity and sophistication of cyber threats. AI, on the other hand, has a strong ability to learn and adapt to new threats, making it a powerful cyber defense tool.
@@
In addition, the autonomous nature of these AI-based threat response systems represents a significant advantage. They are capable of continuously monitoring network activity, proactively identifying potential vulnerabilities, and promptly responding to mitigate risks [57]. By automating threat detection and response processes, organizations can greatly enhance their security posture and reduce the time gap between threat identification and resolution. Once the benefits of AI have been explored, it is also important to recognize that this technology comes with certain limitations and risks. The study conducted in [58] discusses some of these disadvantages. Among these challenges, one involves the inherent unpredictability of AI, which concerns the possibility that AI responses or actions may deviate from the intended goals or objectives. Furthermore, authors highlight that the creation of a super-intelligent AI could result in more harm than benefit when exploited by malicious entities. This underlines the urgency of managing the unpredictability that is inherent in AI systems. Another study [59] identifies additional negative applications that can be obtained by AI exposing potential risks and ethical concerns. For example, AI can be used as a tool to generate deep fakes [60], which can serve as a powerful disinformation weapon used by companies or states, capable of impacting vast numbers of individuals due to the widespread reach of the Internet. Another use can be for password-guessing tools [61] or the impersonation of humans on social media platforms [62]. Through sophisticated algorithms, AI can replicate the behavior of real individuals, seeking personal advantages or even financial gain. The last of the malicious tools proposed in the study is the use of AI for Supported Hacks, where AI is trained to execute various attacks, in this case, Wi-Fi attacks, through the Pwnagotchi 1.0.0 tool [63], to deauthenticate users from wireless networks.
@@
When integrating AI into threat response systems, there are various goals to accomplish. First, is to achieve an autonomous selection of measures to counter ongoing cyber threats. A second goal is the ability that involves the restoration of the system to a stable state. Therefore, these systems must have predefined guidelines to prevent and restore from diverse identified attacks. The preventive defense measures are called countermeasures. Although it is important to note that there can also be reactive countermeasures and not only preventive countermeasures. Countermeasures can be defined as the actions that must be taken to effectively stop an attempted attack. In the field of computing and telecommunications, examples of countermeasures are “shutting down specific ports”, “cutting off external communications”, “deleting files” or “terminating running processes”, etc.
@@ III Features
It is essential to remark that these chosen features constitute one potential proposal among various alternatives. Other features could potentially be gathered and analyzed, but we firmly consider that the selected are the most advantageous choices. This conviction is based in the fact that these features represent in a simple and graphic way the instruments or methodologies used in the proposal of each of the selected papers. With these features, a reader can gain an overview of the most common approaches used in the world of reaction systems, which scenarios are most likely to be addressed, and even the rationale behind the algorithms for the search for countermeasures.
In the following sections, each of the features is described, explaining each of the values that have been considered most relevant in our perspective. The rest of the values are listed in section III-F and are further explained in section IV.
@@ III-A Methodology
  • ML: An evolving discipline in which computer algorithms are designed to emulate human intelligence through computer-based learning based on a digital environment [69]. From this digitalized context, an ML approach could be used to identify and learn from patterns within the collected data. In particular, it could be used in a reactive framework for example to dynamically adjust countermeasures according to the latest attack trends and tactics, or optimize resource allocation by calculating the costs of different countermeasures.
  • Bayesian Networks: These are specific types of graphical models. The specific type are Directed Acyclic Graphs (DAG) [70, 71]. In other words, all the edges of the graph are directed and there are no cycles. Thus, to solve the tasks, probabilistic relations are used as a working model on these graphs. Specifically, it could be used for risk calculations against the selection of countermeasures in a reactive environment. Figure 2 shows an example of a Bayesian network. It shows the probabilities for each branch and the different nodes of the network representing different states.
    Figure 2: Graphical Bayesian Network example
  • Attack Graphs: Graphical description used in cybersecurity to represent and visualize potential attack paths and sequences that an adversary could follow to compromise a computer system or network [48]. It can be used to analyze vulnerabilities and countermeasures, among other possibilities. Figure 3 graphically depicts an example attack represented through an attack graph.
    Figure 3: Attack Graph example

### v2

@@ Abstract
There is no denying that the use of Information Technology (IT) is undergoing exponential growth in today’s world. This digital transformation has also given rise to a multitude of security challenges, notably in the realm of cybercrime. In response to these growing threats, public and private sectors have prioritized the strengthening of IT security measures. In light of the growing security concern, Artificial Intelligence (AI) has gained prominence within the cybersecurity landscape. This paper presents a comprehensive survey of recent advancements in AI-driven threat response systems. To the best of our knowledge, the most recent survey covering the AI reaction domain was conducted in 2017. Since then, considerable literature has been published, and therefore, it is worth reviewing it. In this comprehensive survey of the state of the art reaction systems, five key features with multiple values have been identified, facilitating a homogeneous comparison between the different works. In addition, through a meticulous methodology of article collection, the 22 most relevant publications in the field have been selected. Then each of these publications has been subjected to a detailed analysis using the features identified, which has allowed for the generation of a comprehensive overview revealing significant relationships between the papers. These relationships are further elaborated in the paper, along with the identification of potential gaps in the literature, which may guide future contributions. A total of seven research challenges have been identified, pointing out these potential gaps and suggesting possible areas of development through concrete proposals.
@@
Artificial Intelligence, Cybersecurity, Reaction, Countermeasure, Survey
@@ I Introduction
The conjunction between AI and cybersecurity reaction systems can bring a powerful combination that amplifies the effectiveness of cybersecurity mechanisms. For example, where IRS focuses on real-time threat monitoring, analysis, and incident response, AI increases these capabilities by taking advantage of advanced analytics, ML, and predictive models. In this sense, the collaboration enables the identification of sophisticated attack patterns, the prediction of unknown future threats, as well as the generation of actionable perspectives to inform decision-making processes [33].
Some researchers have recognized the significance of incorporating AI techniques into cybersecurity mechanisms [34]. Furthermore, numerous papers have proposed a variety of approaches to develop more effective countermeasure selection mechanisms that use AI. Therefore, it is crucial to review and extend the findings obtained from the studies by integrating the latest advances in AI and capturing the knowledge gained from subsequent research efforts. This is why it is important to review the state of the art of threat response systems using AI so that all developments or approaches in the field can be brought into a unified vision.
@@
There have been previous studies [37, 38, 39, 40, 41, 42] that unify previously conducted work investigating the automaticity of countermeasure selection and reaction strategies. However, due to the rapid pace of advancements in the field of reaction mechanisms using AI, there is a need for an up-to-date study to consolidate existing knowledge, establish benchmarks, and compare different modern approaches. As far as we are aware, the latest survey addressing the field of AI reaction was carried out in 2017. Therefore, by conducting a comprehensive study, a holistic understanding of the current state of the art can be gained, research gaps can be identified, and valuable information can be provided for researchers, practitioners, and organizations working to improve cybersecurity defenses through AI-based reaction systems. In addition to the above, it is important to emphasize that this paper extends beyond a narrow focus on threat response. Instead, it aims to explore the convergence of cybersecurity and AI, reflecting on their current state of integration. Consequently, our analysis encompasses not only IRS but also evaluates works featuring AI-driven reactive mechanisms such as algorithms, frameworks, and architectures. By addressing both challenges and advancements in this domain, our study illuminates the complexities and opportunities inherent in safeguarding systems from cyber threats.
In order to carry out this study, a recent state-of-the-art recollection work has been carried out, following the methodology described in Figure 1. The diagram illustrates two types of selection: objective and subjective. In particular, the research started with a literature search based on inclusive criteria, focusing on the intersection of cyber defense and AI between 2018 and 2023. This criterion was applied in reputable search engines such as Google Scholar, Science Direct, IEEE Xplore, and Research Gate, with additional keywords such as “reaction”, “countermeasure”, “cyber defense”, “AI”, and “autonomous countermeasure system”. During the initial selection phase, materials not directly published in journals or conferences were excluded. Preference was given to publications with a higher impact factor to ensure the inclusion of high-quality academic work.
After objective filtering, the abstracts of the pre-selected papers were reviewed to identify those most aligned with the research objectives. This subjective assessment took into account factors such as relevance, novelty, and methodological rigor. Subsequently, each selected article was subjected to an exhaustive reading that allowed us to identify the most outstanding articles of the last five years. In total, 22 exemplary articles were selected.
Our research revealed a critical gap in current field studies, i.e., the limited exploration of AI-based solutions for reactive countermeasures. Our analysis of the state of the art revealed significant underdevelopment in the field of reactive defense strategies, leaving a heavy reliance on human intervention amongst the increasing complexity of modern cyber threats. Thus, the advent of AI offers an important opportunity that is crucial to study and expand.
Consequently, this study was conducted to fill such a gap and shed light on the current state of such systems to ultimately identify their main challenges, trying to unify the research found so far under the same common characteristics and to create a comprehensive picture. By analyzing existing work and outlining future challenges, the paper aims to motivate the research community to further explore and develop AI-integrated reactive strategies in cybersecurity. Therefore, the paper aims to contribute significantly to the field by providing the following:
  1. 1.
    Related studies are identified and synthesized to enrich community knowledge, providing guidance based on the strengths and weaknesses of the selected papers.
  2. 2.
    Side-by-side comparisons to elucidate key trends and emerging challenges within the domain of AI-integrated cybersecurity strategies.
  3. 3.
    Outlining potential research directions to stimulate further inquiry and innovation in this critical area of study.
@@
Figure 1: Methodology used for the selection of the papers to be analyzed in search of a comprehensive overview of the state of the art.
@@ II Background
For the sake of the reader, the main objective of this section is to cover several key concepts that form the basis of AI-powered threat response systems. Particularly, understanding these concepts is essential to understanding the context and importance of this research ecosystem. As discussed previously, the increase in the number of cyber-attacks is tangible. This is why systems such as IDS, which are dedicated to detecting threats, are used. And, it is because of the inability of IDS to react that they are combined with response systems such as IPS or IRS.
Previously, cyber-attack detection and prevention systems heavily relied on static techniques such as rule-based analysis [43] or approaches built upon hash-based analysis [44] or the normal behavior of applications [45]. However, with the advancement of technology and the increasing sophistication of cyber threats, static detection has become inefficient [46]. Therefore, there is a growing need for more advanced and dynamic approaches to effectively safeguard against evolving attack vectors, that can cover modern attacks.
@@
For a better comprehension understanding of an AI-driven response system, in conjunction with other essential components of a network (IDS, IPS), Figure 2 has been designed. In this figure, the different components are graphically displayed. The logical flow is divided in steps, that correspond to the numbers associated with the figure:
  1. 1.
    The scenario describes an attack from a malicious actor where a device is infected.
  2. 2.
    After infection, the IDS would be in charge of detecting the malware on the infected device. AI is particularly useful at this stage, as it can analyze vast amounts of data in real time, recognizing patterns and anomalies that might indicate an intrusion.
  3. 3.
    Once identified the malware, the IDS would send an alarm to the IRS. The IRS then analyzes the nature of the threat and selects the most suitable and effective countermeasure. AI’s role here is crucial, as it can analyze historical data and current threats to suggest optimal responses.
  4. 4.
    After analyzing the alert it receives from the IPS, the IRS would then select the most effective countermeasure and implement it.
  5. 5.
    Meanwhile, an IPS is constantly analyzing the network and proactively generating security measures. The AI-powered IPS acts as a proactive barrier, preventing potential intrusions before they can exploit vulnerabilities in the network.
Figure 2: Prevention, response, and detection AI-based systems scenario.
Technologies such as AI can offer a new approach that leverages the latest advancements to create highly efficient systems[50, 51]. These new developments have the potential to enhance on many levels, for instance in the fields of medicine [52], robotics [53], finance [54], and transport [55], among others.
@@
For instance, one main advantage of AI in cybersecurity is its ability to analyze large volumes of data and identify patterns and anomalies that may be indicative of a cyber-attack [56]. For example, algorithms using ML can be trained on large amounts of datasets to recognize different patterns of known attacks and detect previously unseen threats, such as “zero-days” vulnerabilities [57]. As already said, traditional approaches to security rely on static, rule-based systems that are unable to handle the increasing complexity and sophistication of cyber threats. AI, on the other hand, has a strong ability to learn and adapt to new threats, making it a powerful cybersecurity tool.
When it comes to cybersecurity, the deployment of these advances can help for example in improving the detection, prevention, and response systems discussed above.
Particularly, AI can potentially enhance IDS by enabling it to analyze extensive datasets, thereby improving detection accuracy and minimizing false positives through behavioral analysis. By analyzing the behavior of users and systems, AI algorithms could learn normal behavior patterns, allowing them to identify deviations or anomalies that suggest possible intrusions or malicious activities, even in encrypted traffic [58]. Additionally, an IDS could be integrated with threat intelligence feeds, continually updating its understanding of emerging threats and attack vectors.
Similarly, IPS systems can benefit from AI by enabling dynamic adjustment of security policies or cybersecurity strategies, informed by evolving threat landscapes and historical data insights. With AI-driven IPS, rapid discrimination between benign network activities and malicious intrusions becomes possible. Real-time analysis of network traffic by AI algorithms enables the dynamic adaptation of firewall rules and access controls to counter emerging threats. Similarly, as in IDS systems, AI can support IPS systems by identifying zero-day threats through analysis of behavioral patterns and anomalies.Integrating AI into IPS systems can also present a significant advantage in handling massive amounts of data. AI algorithms excel at processing and analyzing large data sets, allowing IPS systems to extract valuable information from the growing volume of network traffic and security logs.
Within the IRS ecosystem, by deploying natural language processing techniques, AI could facilitate the rapid understanding and categorization of incident reports, accelerating the prioritization and allocation of resources for response efforts. AI-driven automation accelerates the execution of pre-defined response procedures, enabling security incidents to be quickly contained and mitigated. In addition, AI-powered IRS could automate the initial triage of security incidents by analyzing incoming alerts and prioritizing them based on severity, impact, and likelihood of exploitation. In addition, AI in an IRS system could perform predictive analytics to anticipate future security incidents by leveraging historical data and current threat intelligence. Another benefit of this integration could be the ability to quickly and efficiently select the optimal set of countermeasures to address threats in real time.
In addition, the autonomous nature of these AI-based threat response systems represents a significant advantage. They are capable of continuously monitoring network activity, proactively identifying potential vulnerabilities, and promptly responding to mitigate risks [59]. By automating threat detection and response processes, organizations can greatly enhance their security posture and reduce the time gap between threat identification and resolution.
@@
Once the benefits of AI have been explored, it is also important to recognize that this technology comes with certain limitations and risks. The study conducted in [60] discusses some of these disadvantages. Among these challenges, one involves the inherent unpredictability of AI, which concerns the possibility that AI responses or actions may deviate from the intended goals or objectives. Furthermore, authors highlight that the creation of a super-intelligent AI could result in more harm than benefit when exploited by malicious entities. This underlines the urgency of managing the unpredictability that is inherent in AI systems. Another study [61] identifies additional negative applications that can be obtained by AI exposing potential risks and ethical concerns. For example, AI can be used as a tool to generate deep fakes [62], which can serve as a powerful disinformation weapon used by companies or states, capable of impacting vast numbers of individuals due to the widespread reach of the Internet. Another use can be for password-guessing tools [63] or the impersonation of humans on social media platforms [64]. Through sophisticated algorithms, AI can replicate the behavior of real individuals, seeking personal advantages or even financial gain. The last of the malicious tools proposed in the study is the use of AI for Supported Hacks, where AI is trained to execute various attacks, in this case, Wi-Fi attacks, through the Pwnagotchi 1.0.0 tool [65], to deauthenticate users from wireless networks.
@@
When integrating AI into threat response systems, such as an autonomous IRS, there are various goals to accomplish. First, is to achieve an autonomous selection of measures to counter ongoing cyber threats. A second goal is the ability that involves the restoration of the system to a stable state. Therefore, these systems must have predefined guidelines to prevent and restore from diverse identified attacks. The preventive defense measures are called countermeasures. Although it is important to note that there can also be reactive countermeasures and not only preventive countermeasures. Countermeasures can be defined as the actions that must be taken to effectively stop an attempted attack. In the field of computing and telecommunications, examples of countermeasures are “shutting down specific ports”, “cutting off external communications”, “deleting files” or “terminating running processes”, etc.
@@ III Features
The selection of features to compare the surveyed papers was a meticulous process based on a structured methodology shown in Figure 3. In particular, we began by conducting an extensive literature review to understand the research landscape relevant to the reaction ecosystem. This review revealed many characteristics that were found in a variety of studies. These features were then subjected to a filtering process aimed towards isolating those consistently present in the papers studied. This initial filtering ensured that only the most prevalent and relevant features were considered. In addition, a further filter was applied to select those characteristics that would facilitate a meaningful comparison in the context of this study.
Figure 3: The methodology employed for selecting the comparative attributes to analyze the selected papers.
It is essential to remark that these chosen features constitute one potential proposal among various alternatives. Other features could potentially be gathered and analyzed, but we firmly consider that the selected are the most advantageous choices. This conviction is based on the fact that these features represent in a simple and graphic way the instruments or methodologies used in the proposal of each of the selected papers. With these features, a reader can gain an overview of the most common approaches used in the world of reaction systems, which scenarios are most likely to be addressed, and even the rationale behind the algorithms for the search for countermeasures.
In the following sections, each of the features is described, explaining each of the values that have been considered most relevant in our perspective. Furthermore, we elaborate on the impact that these features could have on responses to cyber-attacks and AI-based threat response systems. The aim is to establish the connections and effects between the identified characteristics and their importance for effectively mitigating cyber threats. Through this approach, a comprehensive understanding of how each characteristic contributes to the overall cyber defense strategy and resilience against evolving threats is provided. The rest of the values are listed in section III-F and are further explained in section IV.
@@ III-A Methodology
  • Machine Learning (ML): An evolving discipline in which computer algorithms are designed to emulate human intelligence through computer-based learning based on a digital environment [71]. From this digitalized context, an ML approach could be used to identify and learn from patterns within the collected data. In particular, in order to counteract cyber-attacks, it could be used in a reactive framework, for example, to dynamically adjust countermeasures according to the latest attack trends and tactics, or optimize resource allocation by calculating the costs of different countermeasures. In addition, ML algorithms can play a crucial role in optimizing the allocation of resources by calculating the costs associated with different countermeasures.
  • Bayesian Networks: These are specific types of graphical models. The specific type are Directed Acyclic Graphs (DAG) [72, 73]. In other words, all the edges of the graph are directed and there are no cycles. Thus, to solve the tasks, probabilistic relations are used as a working model on these graphs. Specifically, it could be used for risk calculations against the selection of countermeasures in a reactive environment. For instance, it could analyze incoming data from security sensors, assessing anomalies and patterns indicative of cyber threats. In addition, they can facilitate the development of effective incident response plans by modeling the dependencies between different response actions and their outcomes. Figure 4 illustrates a Bayesian network designed to model various states and probabilities associated with the failure of a system, either due to malware exploitability or hardware failure. The network consists of several nodes representing different states within a security incident. Each branch of the network signifies a potential outcome, while the nodes encapsulate the probability of these outcomes. Through this representation, one can discern the potential scenarios that may unfold during a security breach, along with their associated probabilities.
    Figure 4: Probabilistic analysis of computer system failures represented through a graphical Bayesian Network.
  • Attack Graphs: Graphical description used in cybersecurity to represent and visualize potential attack paths and sequences that an adversary could follow to compromise a computer system or network [49]. Attack graphs are a versatile tool, capable of analyzing vulnerabilities and countermeasures. Thereby, these graphs can help to identify interconnected vulnerabilities and their impact on the overall security posture. Consequently, organizations can develop accurate and tailored mitigation strategies to effectively address these weaknesses. Figure 5 graphically depicts an example attack represented through an attack graph. In the graph, an example of how the attack could start via a browser can be seen. Through spear phishing a user account, and through the execution of the attack, the computer could be infected, reaching the operating system itself.
    Figure 5: Graphical representation of an attack against an operating system using an Attack Graph.